OpenStudy (sh3lsh):

Find double integral of 6xy+6ydA with respect to the bounds x^2+y^2=16,y=0,y=3x.

OpenStudy (sh3lsh):

My work: \[\int\limits_{0}^{4}\int\limits_{0}^{3x}6xy+6y dy dx\] = 2034 Unfortunately, it's incorrect

OpenStudy (sh3lsh):

=**2304

OpenStudy (perl):

did you graph the region over which we are integrating

OpenStudy (perl):

lets do that first

OpenStudy (sh3lsh):



OpenStudy (perl):

one way you can do it

OpenStudy (perl):



OpenStudy (sh3lsh):

Polar? I thought the calculation wouldn't be fun. Because the function is 6xy + 6y

OpenStudy (perl):

i think you can do it rectangularly

OpenStudy (perl):

you have to find that intersection point `c`, where does x^2+y^2 = 16 intersect y = 3x, that x value

OpenStudy (perl):

(2/5)*sqrt(10)

OpenStudy (sh3lsh):

Oh! That makes a lot of sense! I see my problem. Thanks so much!! Would rectangularly be the best way to solve it by the way?